OpenStudy (anonymous):

will award medal Find an equation of the tangent plane to the given surface at the specified point. z = sqrt(xy)

OpenStudy (anonymous):

p=(3,3,3)

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

so i keep getting the answer of z=2x+2y-9, from the partial of y and the partial of x evaluated at 3,3,3. and using the fact that z-z_0=f_x(x_o,y_o)(x-x_o)+f_y(x_o,y_o)(y-y_o)

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

but my online says its worng

OpenStudy (anonymous):

You have \[f(x,y)=\sqrt{xy}~~\implies~~\nabla f(x,y)=\left(\frac{\sqrt y}{2\sqrt x},\,\frac{\sqrt x}{2\sqrt y}\right)\] So at \((x,y)=(3,3)\), you get \(\nabla f(3,3)=\left(\dfrac{1}{2},\dfrac{1}{2}\right)\). The tangent plane would then be \[z-3=\frac{1}{2}(x-3)+\frac{1}{2}(y-3)\]

OpenStudy (anonymous):

sorry that coding is really confusing

OpenStudy (anonymous):

Is the latex not showing up?

OpenStudy (anonymous):

no not for me

OpenStudy (anonymous):

oh....

OpenStudy (anonymous):

do its not regular differentiating partially

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

im just so confused on the partial derivitive. my teacher must have not clarified what it means

OpenStudy (anonymous):

oh pellet so would you use the product rule and seperate

OpenStudy (anonymous):

@mathmate l

OpenStudy (mathmate):

In case you were not clear about: \(\nabla f(x,y)\) It means \(\nabla f(x,y)=\left(\dfrac{\partial f}{\partial x},\dfrac{\partial f}{\partial x}\right )=\left(\dfrac{\sqrt y}{2\sqrt x},\,\dfrac{\sqrt x}{2\sqrt y}\right)\)